OpenStudy (anonymous):

The information given represents lengths of sides of a right triangle with c the hypotenuse. Find the correct missing length to the nearest hundredth. A calculator may be helpful. a = ?, b = 13, c = 21 A. 15.83 B. 15.87 C. 16.19 D. 16.49

OpenStudy (anonymous):

You could use the Pythagorean Theorem.

OpenStudy (anonymous):

ive tried and i dont get it at all

OpenStudy (anonymous):

They gave you b and c. The pythagorean theorem is \[a^2+b^2=c^2\] Given that, can you put the numbers in and solve for it?

OpenStudy (anonymous):

would the answer be D?

OpenStudy (anonymous):

\[a^2+13^2=21^2\] \[a^2+169=441\] \[a^2=441-169\] \[a^2=272\] \[\sqrt{a^2}=\sqrt{272}\] \[a=16.4924225025\] Which is rounded to 16.49 so yes, D.

OpenStudy (anonymous):

thank you. i understand it now.